# 22footnotetext: Corresponding author

### v1

@@ 3.4.4. Multimodal Learning
Xie et al. (Xie et al., 2024a) introduced a U-Net-like architecture called Fusion Mamba which is designed for encoding multimodal images and then decoding them. Fusion Mamba shown in Figure 19 fuses features from two different source images. The encoder part incorporates Dynamic Vision State Space (DVSS) block which integrates Efficient State Space Module. ESSM uses an Efficient 2D Selective Scan (ES2D) and Efficient Channel Attention (ECA) (Wang et al., 2020). ECA involves adaptive average pooling followed by a 1D convolution layer and a sigmoid activation function. ECA enables SSM to learn channel representations effectively. The output from DVSS is residually added to the input through a Learnable Descriptive Convolution (LDC). LDC uses a learnable matrix ($m$) combined with an identity matrix. A dot product is performed between the modified learnable matrix and the input map. The result is multiplied in the next step to produce the output map which allows SSM to capture the textural features present in each modality. Dynamic Feature Fusion Module (DFFM) is used to fuse features between different modalities. The DFFM includes Dynamic Feature Enhancement Module (DFEM) for each modality which takes two different feature modes as input which is followed by performing coarse-grained fusion and then passing output through the Cross Modality Fusion Mamba Module (CMFM). The Cross Modality Fusion Mamba (CMFM) block starts by applying layer normalization, followed by a linear layer and depthwise convolution for each modality. The resulting representations from each modality are multiplied and then residually connected with outputs of depth-wise convolutions to resemble a gating mechanism. These representations are further processed through an Efficient Selective 2D Scan (ES2D). The outputs from layer normalization are then passed through a linear layer where it is multiplied with representations from ES2D and residually added. Finally, merged features are passed through another linear layer, followed by the ECA. The decoder contains a patch-expanding block followed by two DVSS blocks. The combined features from DFFM act as skip connections for the decoder. The patch-expanding block can either use transposed convolution or bilinear upsampling of features. Finally, the fused image is obtained from each image modality.
Zhou et al. (Zhou et al., 2024) proposed MGI which is a multimodal approach used for aligning images and gene modalities using a pre-training approach similar to CLIP (Radford et al., 2021). Mamba-based encoder is employed for both images and genes. Contrastive loss is applied between modality embeddings to generate a matrix similar to CLIP. For task-specific learning, the model integrates an attention integration module and a mask output module. In the attention integration module, alignment between modules is achieved through a self-attention block on the gene modality, gene-to-image attention block, MLP layer and then an image-to-gene attention block to combine modalities. Similarly, in the mask output module a gene-to-image attention block is followed by MLP and a dot product is applied between image and gene modalities to generate a mask image which is trained using Dice loss.
Fang et al. introduced GFE-Mamba (et al., 2024d), a model employing a multi-stage training strategy. First, they train a 3D GAN to convert MRI images into PET images. Afterward, latent representations from MRI and PET are concatenated with tabular data, where continuous features pass through a linear layer, and discrete features pass through an embedding layer, before being fed into a Mamba classifier. The representations from Mamba classifier and MRI and PET latents from 3D GAN undergo a pixel-level bi-cross attention operation, applying attention operation between Mamba classifier’s representation and MRI/PET latents respectively. GFE-Mamba integrates a U-Net-based architecture, with the encoder in the 3D U-Net utilizing three downsample blocks, each composed of a convolution block, ReLU activation, max pooling, and group normalization, with encoder channel sizes of 64, 128, and 256. The middle layer features ViT blocks to enhance representation capture, outperforming traditional residual layers used in conventional 3D GANs. The decoder consists of upsample blocks, each with a transposed convolution and ReLU activation, with channel sizes of 256, 128, and 64. The 3D GAN is trained using a discriminator, similar to the standard GAN training method with real and fake images. The pixel-level bi-cross attention combines cross-attention from MRI and PET latent spaces with Mamba representations. First, attention between MRI latents and Mamba is calculated with RMS normalization, followed by attention from the PET block, again with RMS normalization, and the final output is classified into binary classes.
3.5. Applications in Various Medical Domains
@@
3.5.1. Medical Image Segmentation
Medical image segmentation is a technique used to identify and extract specific Regions Of Interest (ROI) from medical images such as tumors, lesions, tissues or organs. The objective is to divide the image into areas that share similar features including color, texture, brightness, and contrast. Table 3 outlines the overview of segmentation models, parameters, descriptions and the code availability. Figure 20 shows the workflow of Mamba-based models in medical image segmentation task. Some notable research works on mamba-based models for medical image segmentation includes:
Figure 20. Work Flow of Segmentation Task
Xing et al. (Xing et al., 2024) proposed an architecture called SegMamba along with a new dataset comprising 500 3D computed tomography scans with expert annotations. The proposed model features an architecture similar to the transformer-based U-Net. Initially, it employs depthwise convolution with a kernel size of $7\times 7\times 7$, padding of $3\times 3\times 3$, and a stride of $2\times 2\times 2$. On the encoder side, the model incorporates a series of TS-Mamba blocks and downsample blocks with residual connections mirrored on the upsample side. TS-Mamba block has a transformer-like architecture that uses gated spatial convolution to capture spatial relationships between 3D features. This is followed by the Tri-orientated Spatial Mamba (TOM) block, which replaces the traditional multi-headed attention layers found in transformers. TOM applies independent Mamba models to forward, backward, and inter-slice feature interactions respectively. The computed values are then processed by a multilayer perceptron to enrich feature representation. Each TS-Mamba block ends with downsampling, doubling the number of channels while halving the height, width, and depth of the features. These outputs are then residually connected to the decoder side of the network. The decoder uses convolution transpose networks to upsample the image. SegMamba was evaluated on three publicly available datasets: the BraTS2023 dataset, AIIB2023 dataset, and CRC-500. It demonstrated superior performance across all metrics compared to other transformer-based U-Net models.
Table 3. Overview of Segmentation Models
@@ 3.5.1. Medical Image Segmentation
SegMamba (Xing et al., 2024) - TSMamba Multiple Tumor analysis
H-Vmunet (Wu et al., 2024a) 8.97 H-VSS Skin lesion and polyp analysis
LKM-UNet (Wang et al., 2024a) - Bi-Mamba 2D & 3D Multiple organ analysis
Mamba-UNet (Wang et al., 2024d) - VMamba Multiple organ analysis
@@
Weak Mamba UNet (Wang and Ma, 2024) - VMamba Analysis of heart with WSL framework
LightM-UNet (Liao et al., 2024) 1.87 RVM Tumor analysis in chest X-ray images
@@
UltraLight-VM UNet (Wu et al., 2024b) 0.049 Mamba Skin lesion analysis with light weight networks
T-Mamba (Hao et al., 2024) 1.04 Tim 2D & 3D tooth analysis
HC-Mamba (Xu, 2024) 13.88 Mamba Skin lesion analysis
@@
Semi-Mamba-UNet (Ma and Wang, 2024) - VMamba Analysis of heart with SSL framework
ViM-UNet (Archit and Pape, 2024) 18 ViM Analysis of microscopic organs(cell)
U-Mamba (Ma et al., 2024) - Mamba 2D & 3D Multiple organ analysis
@@
Mamba-HUNet (Sanjid et al., 2024) - VMamba Analysis of Multiple Sclerosis Legion in brain
UU-Mamba (Tsai et al., 2024) - Mamba Heart Analysis integrated with SAM optimizer
CAMS-Net (Khan et al., 2024) - NC-Mamba Heart region segmentation
MUCM-Net (Yuan et al., 2024) 0.047 Mamba Skin lesion analysis
@@
Wu et al. (Wu et al., 2024a) introduced High-order Vision Mamba UNet (H-Vmunet), which is also similar to a transformer-based U-Net architecture. This work addresses challenges in skin lesion analysis (ISIC2017), spleen cancer detection, and polyp detection (CVC-ClinicDB dataset). The authors introduced H-VSS module, a modified version of VSS, replacing it with the H-SS2D module. In H-SS2D module, the input is projected into $2C$ channels. The split of these channels depends on the order of the H-SS2D module. H1-SS2D splits channels by half, while H3-SS2D splits them into $C/4$ and $47C/$ 4 for SS2D and Local-SS2D, respectively. Local-SS2D, a submodule of H-VSS, applies convolution 2D and SS2D operations to each half of channels. Each step in H-VSS involves multiplication from the projection to Local-SS2D and scanning with SS2D layer. The number of scans increases with the order of the selective scan. Additionally, SS2D employs scan expansion, performed in four different scan order directions: top-left to bottom-right, top-right to bottom-left, bottom-left to top-right and bottom-right to top-left. SS2D uses S6 SSM as described in Mamba. The overall model architecture is based on U-Net, starting with two convolutional layers, followed by H-VSS modules of order 2, 3, 4, and 5 from layers 3 to 6, each followed by a convolutional layer with a kernel size of 3. The residual layers of this U-Net uses shared weights of Channel Attention Bridge (CAB) and Spatial Attention Bridge (SAB) for information fusion between the encoder and decoder. On the decoder side, the model starts with H-VSS blocks of order 5, 4, 3, and 2 from layers 1 to 3, with the addition of features from SAB and CAB. The features are upsampled using bilinear interpolation. H-Vmunet demonstrates superior performance across metrics in the ISIC2017, Spleen, and CVC-ClinicDB datasets.
Wang et al. (Wang et al., 2024a) devised Large Kernel Vision Mamba UNet (LKM-UNet) for both 2D and 3D segmentation tasks. The encoder of LKM-UNet uses depthwise convolution which is followed by LM block and a downsampling block. LM block consists of pixel-wise SSM and patch-level SSM, each followed by a bidirectional mamba block at each SSM level. The bidirectional Mamba block includes forward and backward SSMs which is similar to ViM. The representations from pixel-level SSMs are processed by bidirectional SSM, which is then pooled, and passed to patch-level SSM, that employs another bidirectional Mamba block. This allows features from each sub-window to interact with patch-level SSMs, followed by a bidirectional Mamba. The pixel-level SSM representations are then residually added to unpooled representations from patch-level SSMs. After LM block, downsampling occurs, doubling the feature space while halving the height, width, and depth. In the decoder, LKM-UNet uses convolution-based upsampling with residual connections from corresponding encoder layers. Figure 21 shows the architecture of LKM-UNet and its components LM Block and bidirectional Mamba. LKM-UNet outperforms traditional UNet and mamba-based methods on 3D multi-organ segmentation with Abdomen CT dataset and 2D segmentation with 110 MRI cases in Abdomen MR dataset.
Figure 21. Architecture of LKM-UNet (Wang et al., 2024a)
@@
Wang et al. (Wang et al., 2024d) presented Mamba-UNet, a model that adapts a purely ViM-based U-Net architecture with linear embedding and VSS blocks. This network includes encoder, decoder, and bottleneck layers. Initially, medical images, defined by height and width, are converted into a 1D sequence. An embedding layer transforms them into a size denoted by C, followed by VSS blocks. During downsampling, the feature space is doubled while height and width are halved. The bottleneck layer consists of two VSS blocks without downsampling. In decoder, features are reconstructed by upsampling, reducing feature space by half and doubling the height and width. Skip connections are employed between corresponding encoder and decoder layers. Mamba-UNet demonstrates superior performance compared to attention-based and transformer-based U-Net models on datasets such as the Automated Cardiac Diagnosis Challenge, which includes MRI scans of 100 patients, and the Synapse Multi-Organ Segmentation Challenge, which contains 30 abdominal CT scans with expert annotations.
Wang et al. (Wang and Ma, 2024) proposed Weak Mamba UNet, which combines CNNs, ViTs, and Visual Mamba for WSL. This model uses a weighted mixture of the three architectures, each having a symmetric architecture with identical input and output sizes. For CNN-based model, the authors utilized classical U-Net with convolution layers for downsampling. ViT-based UNet and Visual mamba-based UNet follow a similar symmetric architecture. Weak Mamba UNet uses scribble annotations from images, and all networks are evaluated using a pseudo-label loss function and a scribble loss. The authors also experimented with using three instances of mamba based UNet, ViT-based UNet, and CNN-based UNet separately. However, the combined approach using CNN, ViT, and mamba-based UNet outperformed the other configurations. Weak Mamba UNet showed superior performance compared to other ViT-based UNet models, such as Swin-UNet, in WSL settings, particularly excelling on ACDC dataset.
Liao et al. (Liao et al., 2024) introduced LightM-UNet, which is an extension of VSS module by incorporating a Residual Vision Mamba (RVM) Layer. RVM includes layer normalization, followed by another layer normalization, and then VSS module, with residual connections and scaling. This layer addresses long-term spatial dependencies. Each encoder block in LightM-UNet contains RVM and downsampling layers. The encoder block representations are used for residual connections in corresponding decoder blocks. The decoder blocks consist of depthwise convolution with residual connections scaled similar to RVM layer, and bilinear transformation is used to restore predictions to original resolution. LightM-UNet achieves segmentation with significantly lower number of parameters compared to transformer-based models, reducing computation cost and parameter count by 21×21$\times$21 × and 116×116$\times$116 ×, respectively. It surpasses state-of-the-art performance on two publicly available datasets, LiTs and Montgomery&Shenzhen. Additionally, the authors demonstrated that LightM-UNet can be extended to 3D, evaluated on the LiTs dataset, and still surpass state-of-the-art performance.
Wu et al. (Wu et al., 2024b) devised UltraLight-VM UNet, which employs Parallel Vision Mamba (PVM). Each PVM consists of $N$independent Mamba layers, with input representations divided into $N$ parts along channel dimension $(C/N)$. The authors provided four configurations with 1, 2, and 4 Mamba layers, respectively. Similar to LightM-UNet, Mamba layers in UltraLight-VM UNet use scaled residual connections. All representations from Mamba layers are concatenated, followed by layer normalization and a projection layer, which allows for customizable downsampling of feature space. Unlike LightM-UNet, which uses VSS, UltraLight-VM UNet employs classical Mamba layers, reducing number of parameters and enhancing performance. The encoder layers consist of convolution layers for stages 1-3 and PVM layers for stages 4-6. Residual connections incorporate Spatial Attention Bridge (SAB) and Channel Attention Bridge (CAB) layers, differing from LightM-UNet. The decoder blocks mirror the encoder structure and use bilinear transformation to restore predictions to the original resolution. UltraLight-VM UNet achieves state-of-the-art performance compared to transformer-based methods for skin lesion segmentation on the ISIC2017, ISIC2018, and PH datasets.
Hao et al. (Hao et al., 2024) presented an architecture called T-Mamba. It is a modification of the DenseVNet architecture, incorporating sequential layers of convolutional networks, batch normalization, and the Tim layer. In T-Mamba, Tim layer flattens the input, applies shared position embedding (similar to sinusoidal position embedding), normalizes representations using layer norm, and passes them to a modified ViM block, which includes forward and backward SSM similar to Vision Mamba. Additionally, Tim has a branch that converts input into Fourier domain, extracts features using learnable weight parameters, implements band-pass filtering, and then applies inverse Fourier transform. The Fourier domain representations are multiplied with max-pooled non-SSM part of Mamba representation. A Tim gated selection network is used to select features from forward SSM, backward SSM, and Fourier domain, respectively. The encoder part, inspired by DenseVNet, connects and stacks each layer repeatedly, followed by downsampling up to three layers. Subsequent encoder layers are upsampled by 2×2$\times$2 ×, 2×2$\times$2 ×, and 4×4$\times$4 ×, respectively, leading to a prediction head for segmentation. T-Mamba achieves state-of-the-art performance on 3D Tooth CBCT dataset, outperforming both convolution-based U-Net and transformer-based U-Net models.
Xu et al. (Xu, 2024) proposed HC-Mamba, which uses HC-SSM block composed of a series of HC-Conv blocks and modified Mamba block using HC-Conv. The input is split into two parts: the first part undergoes a series of HC-Conv blocks, where HC-Conv involves a two-step convolution process with dilated convolution followed by depth-wise separable convolution. The second part uses HC-Convolution block instead of the traditional convolution block in Mamba. The representations from both parts are concatenated and shuffled randomly. HC-Mamba employs a patch embedding layer similar to ViT and applies a series of HC-SSM blocks with patch merging. Patch merging doubles feature space and halves height and width of representations at each layer iteration in the encoder. Residual connections are applied in the decoder, originating from patch embeddings. The decoder consists of HC-Mamba blocks with patch merging until the final projection. HC-Mamba outperforms convolution-based U-Net, transformer-based U-Net, and mamba based U-Net models in skin lesion segmentation on ISIC2017 and ISIC2018 datasets, as well as in multi-organ segmentation on Synapse dataset.
@@
Ma et al. (Ma and Wang, 2024) introduced Semi-Mamba-UNet, which leverages self-supervised learning methodology by implementing two networks: Mamba-UNet and UNet. Mamba-UNet incorporates Visual Mamba blocks similar to UNet. In a semi-supervised learning setting, Semi-Mamba-UNet employs three combinations of loss functions: contrastive Loss, semi-supervised loss, and supervised loss. Pixel-level contrastive loss is calculated using a projection layer in the network’s final representations. The loss is determined by taking the $L2$ norm of representations on both labelled and unlabeled data, normalized by the number of input data ($N$). Semi-supervised loss is applied to the network’s final layers using cross-entropy loss and dice loss, with pseudo labels predicted by the network. Supervised loss uses same loss functions as semi-supervised loss but only with labelled data. The authors experimented with different combinations of U-Net and transformer-based models, finding that mamba-based models use fewer parameters and offer better performance. Semi-Mamba-UNet achieves state-of-the-art performance on ACDC and MICCAI datasets. Compared to other attention-based SSL frameworks, they found that using 10% of labeled data is optimal.
Archit et al. (Archit and Pape, 2024) utilized Vision Mamba for cell structure segmentation, achieving superior performance compared to both transformer-based and Convolution-based approaches on U-Net framework. However, the ViM-UNet is not a modified version of existing Vision Mamba architectures but rather implements Vision Mamba along with corresponding downsample and upsample layers to create a U-Net architecture. The authors demonstrated superior performance on datasets such as LIVECell and CREMI. The paper proposed two variations of ViM-UNet: one with a smaller size, termed “Tiny” with 18 million parameters, and another with a larger size, termed “Large” with 39 million parameters.
Ma et al. (Ma et al., 2024) proposes a network called U-Mamba for 2D and 3D biomedical image segmentation. It features an novel hybrid CNN-SSM architecture that effectively captures localized fine-grained features and long-range dependencies in images. UMamba encoder is composed of building blocks, consisting of two successive residual blocks, followed by SSM Mamba, while decoder is composed of residual blocks with skip connections. U-Mamba incorporates self-configuring feature from nnU-Net, and number of network blocks is automatically determined based on the datasets. Additionally, the authors propose two variants of UMamba, namely, “U-Mamba_Bot,” which integrates U-Mamba block only in the bottleneck, and “U-Mamba_Enc,” which utilizes Mamba blocks across all encoder layers. These architectures have been evaluated across four diverse datasets: Abdomen CT, Abdomen MRI, Endoscopy Images, and Microscopy Images. Figure 22 illustrates the architecture of U-mamba and its components.
Figure 22. Architecture of U-Mamba (Ma et al., 2024)
Mamba-HUNet proposed by Sanjid et al. (Sanjid et al., 2024) is a novel architecture that has been designed for reliable and effective medical picture segmentation. Combining the long-range dependency modeling capabilities of SSMs with local feature extraction skills of CNNs, this model combines the best aspects of both Mamba-UNet and a simplified version of the Hierarchical Upsampling Network (HUNet). The design greatly improves processing efficiency by mutating HUNet into Mamba-HUNet and optimizing it into a more efficient variety without sacrificing performance. To enable efficient processing, input grayscale pictures are divided into patches and converted into $1D$ sequences, a method influenced by ViT and Mamba. Mamba-HUNet preserves spatial integrity while extracting hierarchical features by using patch merging layers and VSS blocks. The effectiveness of Mamba-HUNet is evaluated through experiments on MRI scans, particularly in multiple sclerosis lesion segmentation.
@@
Tsai et al. (Tsai et al., 2024) introduced UU-Mamba, a segmentation model that incorporates an uncertainty-aware loss function. This loss function is a combination of three components: dice coefficient loss (region-based), cross-entropy loss (distribution-based), and focal loss (pixel-based), all combined with a learnable parameter sigma. To address the issue of narrow minima in the uncertainty loss, SAM optimization introduces a hyperparameter epsilon to ensure the minima are flatter, thus preventing the network from overly fitting complicated representations. U-Mamba utilizes both UNet-like architecture and Mamba blocks, enabling parameter-efficient computations and achieving superior performance. Each U-Mamba block includes two residual connections: one with a Mamba block followed by a convolution block and Instance Normalization (IN). Skip connections link features from encoder to decoder. UU-Mamba demonstrates superior performance on ACDC dataset compared to transformer-based UNet models.

### v3

@@ 3.4.4 Multimodal Learning
Xie et al. [76] introduced a U-Net-like architecture called Fusion Mamba which is designed for encoding multimodal images and then decoding them. Fusion Mamba fuses features from two different source images. The encoder part incorporates Dynamic Vision State Space (DVSS) block which contains Efficient State Space Module (ESSM). ESSM uses an Efficient 2D Selective Scan (ES2D), Learnable Descriptive Convolution (LDC) and Efficient Channel Attention (ECA) [150]. Dynamic Feature Fusion Module (DFFM) in the encoder is used to fuse features between different modalities. The decoder contains a patch-expanding block followed by two DVSS blocks. The combined features of DFFM act as skip connections for the decoder. The patch-expanding block can either use transposed convolution or bilinear upsampling of features. Finally, the fused image is obtained from each image modality. Zhou et al. [91] proposed MGI which is a multimodal approach used for aligning images and gene modalities using a pre-training approach similar to CLIP [151]. Mamba-based encoder is employed for both images and genes. Contrastive loss is applied between modality embeddings to generate a matrix similar to CLIP. For task-specific learning, the model integrates an attention integration module and a mask output module. Fang et al. introduced GFE-Mamba [92], a model employing a multi-stage training strategy. First, they train a 3D GAN to convert MRI images to PET images. Afterward, latent representations from MRI and PET are concatenated with tabular data, where continuous features pass through a linear layer, and discrete features pass through an embedding layer, before being fed into a Mamba classifier. The representations from Mamba classifier and MRI and PET latents from 3D GAN undergo a pixel-level bi-cross attention operation, applying attention operation between Mamba classifier’s representation and MRI/PET latents respectively.
From this Section 3.4.4, models with Mamba-based designs have shown clear strengths in combining different types of data such as medical images, genetic information and clinical records. Their ability to capture detailed and long-range patterns through components such as DVSS, CMFM and ECA allows for effective fusion of features, making them well suited for tasks such as image segmentation, disease classification and aligning multiple data sources, while keeping computational demands low compared to attention-based models [152, 153, 154, 155]. However, these models tend to be complex, often requiring multi-step training processes and significant effort to set up. In some cases, they still need considerable computing power and their reliance on large, well-annotated datasets can be a barrier in real-world medical environments where such data are not always available.
3.5 Applications in Various Medical Domains
@@
3.5.1 Medical Image Segmentation
Medical image segmentation is a technique used to identify and extract specific Regions Of Interest (ROI) from medical images such as tumors, lesions, tissues or organs. The objective is to divide the image into areas that share similar features including color, texture, brightness, and contrast. Table 3 outlines the overview of segmentation models, parameters, descriptions and the code availability. Figure 16 shows the workflow of Mamba-based models in the medical image segmentation task. Some notable research works on mamba-based models for medical image segmentation include:
Figure 16: Work Flow of Segmentation Task
Xing et al. [16] proposed an architecture called SegMamba along with a new dataset comprising 500 3D computed tomography scans with expert annotations. The proposed model features an architecture similar to the transformer-based U-Net. Initially, it employs depthwise convolution with a kernel size of $7\times 7\times 7$, padding of $3\times 3\times 3$, and a stride of $2\times 2\times 2$. On the encoder side, the model incorporates a series of TS-Mamba blocks and downsample blocks with residual connections mirrored on the upsample side.
Table 3: Overview of Segmentation Models
@@ 3.5.1 Medical Image Segmentation
SegMamba [16] - TSMamba Multiple Tumor analysis
H-Vmunet [17] 8.97 H-VSS Skin lesion and polyp analysis
LKM-UNet [18] - Bi-Mamba 2D & 3D Multiple organ analysis
Mamba-UNet [78] - VMamba Multiple organ analysis
@@
Weak Mamba UNet [79] - VMamba Analysis of heart with WSL framework
LightM-UNet [80] 1.87 RVM Tumor analysis in chest X-ray images
@@
UltraLight-VM UNet [81] 0.049 Mamba Skin lesion analysis with light weight networks
T-Mamba [82] 1.04 Tim 2D & 3D tooth analysis
HC-Mamba [83] 13.88 Mamba Skin lesion analysis
@@
Semi-Mamba-UNet [84] - VMamba Analysis of heart with SSL framework
ViM-UNet [85] 18 ViM Analysis of microscopic organs(cell)
U-Mamba [86] - Mamba 2D & 3D Multiple organ analysis
@@
Mamba-HUNet [87] - VMamba Analysis of Multiple Sclerosis Legion in brain
UU-Mamba [88] - Mamba Heart Analysis integrated with SAM optimizer
CAMS-Net [89] - NC-Mamba Heart region segmentation
MUCM-Net [90] 0.047 Mamba Skin lesion analysis
@@
Wu et al. [17] introduced High-order Vision Mamba UNet (H-Vmunet), which is also similar to a transformer-based U-Net architecture. This work addresses challenges in skin lesion analysis (ISIC2017), spleen cancer detection, and polyp detection (CVC-ClinicDB dataset). The authors introduced H-VSS module, a modified version of VSS, replacing it with the H-SS2D module. In the H-SS2D module, the input is projected into $2C$ channels. The split of these channels depends on the order of the H-SS2D module. H1-SS2D splits the channels by half, while H3-SS2D splits them into $C/4$ and $7C/4$ for SS2D and Local-SS2D, respectively. The overall model architecture is based on U-Net, starting with two convolutional layers, followed by H-VSS modules of order 2, 3, 4, and 5 from layers 3 to 6, each followed by a convolutional layer with a kernel size of 3. The residual layers of this U-Net uses shared weights of the Channel Attention Bridge (CAB) and Spatial Attention Bridge (SAB) for information fusion between the encoder and decoder. On the decoder side, the model starts with H-VSS blocks of order 5, 4, 3, and 2 from layers 1 to 3, with the addition of features from SAB and CAB. The features are upsampled using bilinear interpolation.
@@
Wang et al. [18] devised Large Kernel Vision Mamba UNet (LKM-UNet) for 2D and 3D segmentation tasks. The encoder of LKM-UNet uses depthwise convolution which is followed by LM block and a downsampling block. LM block consists of pixel-wise SSM and patch-level SSM, each followed by a bidirectional mamba block at each SSM level. The bidirectional Mamba block includes forward and backward SSMs which is similar to ViM. In the decoder, LKM-UNet uses convolution-based upsampling with residual connections from the corresponding encoder layers. Wang et al. [78] presented Mamba-UNet, a model that adapts a purely ViM-based U-Net architecture with linear embedding and VSS blocks. This network includes encoder, decoder, and bottleneck layers. Initially, medical images, defined by height and width, are converted into a 1D sequence. An embedding layer transforms them into a size denoted by C, followed by VSS blocks. In decoder, features are reconstructed by upsampling, reducing feature space by half and doubling the height and width. Skip connections are employed between the corresponding encoder and decoder layers. Wang et al. [79] proposed Weak Mamba UNet, which combines CNNs, ViTs, and Visual Mamba for WSL. This model uses a weighted mixture of the three architectures, each having a symmetric architecture with identical input and output sizes. For CNN-based model, the authors utilized classical U-Net with convolution layers for downsampling. ViT-based UNet and Visual mamba-based UNet follow a similar symmetric architecture. Weak Mamba UNet uses scribble annotations from images, and all networks are evaluated using a pseudo-label loss function and a scribble loss.
Liao et al. [80] introduced LightM-UNet, which is an extension of the VSS module incorporating a Residual Vision Mamba (RVM) Layer. RVM includes layer normalization, followed by another layer normalization, and then the VSS module, with residual connections and scaling. This layer addresses long-term spatial dependencies. Each encoder block in LightM-UNet contains RVM and downsampling layers. The decoder blocks consist of depthwise convolution with residual connections scaled similarly to RVM layer, and bilinear transformation is used to restore the predictions to the original resolution. Wu et al. [81] devised UltraLight-VM UNet, which employs Parallel Vision Mamba (PVM). Each PVM consists of $N$ independent Mamba layers, with input representations divided into $N$ parts along the channel dimension $(C/N)$. The authors provided four configurations with 1, 2, and 4 Mamba layers, respectively. Similarly to LightM-UNet, Mamba layers in UltraLight-VM UNet use scaled residual connections. All representations from Mamba layers are concatenated, followed by layer normalization and a projection layer, which allows for customizable downsampling of feature space. The encoder layers consist of convolution layers for stages 1-3 and PVM layers for stages 4-6. Residual connections incorporate Spatial Attention Bridge (SAB) and Channel Attention Bridge (CAB) layers, differing from LightM-UNet. The decoder blocks mirror the encoder structure and use bilinear transformation to restore predictions to the original resolution. Hao et al. [82] presented an architecture called T-Mamba. It is a modification of the DenseVNet architecture, incorporating sequential layers of convolutional networks, batch normalization, and the Tim layer. In T-Mamba, Tim layer flattens the input, applies shared position embedding (similar to sinusoidal position embedding), normalizes representations using layer norm, and passes them to a modified ViM block, which includes forward and backward SSM similar to Vision Mamba. The encoder part, inspired by DenseVNet, connects and stacks each layer repeatedly, followed by downsampling up to three layers. Subsequent encoder layers are upsampled by $\times, 2×2\times$, and $4\times$, respectively, leading to a prediction head for segmentation. Xu et al. [83] proposed HC-Mamba, which uses HC-SSM block composed of a series of HC-Conv blocks and modified Mamba block using HC-Conv. The input is split into two parts: the first part undergoes a series of HC-Conv blocks, where HC-Conv involves a two-step convolution process with dilated convolution followed by depth-wise separable convolution. The second part uses HC-Convolution block instead of the traditional convolution block in Mamba. The representations from both parts are concatenated and shuffled randomly. The decoder consists of HC-Mamba blocks with patch merging until the final projection.
@@
Ma et al. [84] introduced Semi-Mamba-UNet, which leverages self-supervised learning methodology by implementing two networks: Mamba-UNet and UNet. In a semi-supervised learning setting, Semi-Mamba-UNet employs three combinations of loss functions: contrastive loss, semi-supervised loss, and supervised loss. Pixel-level contrastive loss is calculated using a projection layer in the network’s final representations. The loss is determined by taking the $L2$ norm of the representations in the labeled and unlabeled data, normalized by the number of input data ($N$). Semi-supervised loss is applied to the final layers using cross-entropy loss and dice loss, with pseudo labels predicted by the network. Supervised loss uses same loss functions but only on the labeled data. Archit et al. [85] utilized Vision Mamba for cell structure segmentation, achieved better performance compared to both transformer and Convolution based approaches on U-Net frameworks. However, the ViM-UNet is not a modified version of a existing Vision Mamba architectures but rather implements Vision Mamba along with corresponding downsample and upsample layers to create a U-Net architecture. The paper proposed two variations of ViM-UNet: one with a smaller size, termed “Tiny” with 18 million parameters, and another with a larger size, termed “Large” with 39 million parameters.
@@
Ma et al. [86] proposes a network called U-Mamba for 2D and 3D biomedical image segmentation. UMamba encoder is composed of building blocks, consisting of two successive residual blocks, followed by SSM Mamba, while decoder is composed of residual blocks with skip connections. U-Mamba incorporates self-configuring feature from nnU-Net, and number of network blocks is automatically determined based on the datasets. Additionally, the authors propose two variants of UMamba, namely, “U-Mamba_Bot,” which integrates U-Mamba block only in the bottleneck, and “U-Mamba_Enc,” which utilizes Mamba blocks across all encoder layers. Mamba-HUNet proposed by Sanjid et al. [87] is a novel architecture that has been designed for reliable and effective medical image segmentation. This design greatly improves processing efficiency by mutating Hierarchical Upsampling Network (HUNet) into Mamba-HUNet and optimizing it into a more efficient variety without sacrificing performance. To enable efficient processing, input grayscale pictures are divided into patches and converted into $1D$ sequences, a method influenced by ViT and Mamba. Tsai et al. [88] introduced UU-Mamba, a segmentation model that incorporates an uncertainty-aware loss function. This loss function is a combination of three components: dice coefficient loss (region-based), cross-entropy loss (distribution-based), and focal loss (pixel-based), all combined with a learnable parameter sigma. To address the issue of narrow minima in the uncertainty loss, SAM optimization introduces a hyperparameter epsilon to ensure the minima are flatter, thus preventing the network from overly fitting complicated representations. Each U-Mamba block includes two residual connections: one with a Mamba block followed by a convolution block and Instance Normalization (IN).